# Improved polynomial rates of memory loss for nonstationary intermittent dynamical systems

### v1

@@ 3.3. Grossmann–Horner maps
We follow the representation in [20]. Let $T:X\to X$ be a map on $X=[-1,1]$ with graph as in Figure 1, having two surjective and symmetric branches on the subintervals $[-1,0]$ and $[0,1]$, such that $T|_{X\setminus\{0\}}\in C^{1}$ and $T|_{X\setminus\{0,-1,1\}}\in C^{2}$, defined locally by
@@
Here, $a,b>0$, $\eta>0$ and $\gamma>1$ are parameters, $U_{0}(\gamma,\eta)$ is a small neighborhood of 00, and $U_{1_{-}}(\gamma,\eta),U_{-1_{+}}(\gamma,\eta)$ are small neighborhoods of $1,-1$, respectively. For $i=1,2$, the derivatives of $T$ are assumed to satisfy
@@
  • $T$ is strictly monotonous and convex on the two intervals $[-1,0]$ and $[0,1]$, and
@@
Fix $0<a_{-}<a_{+}<\infty$, $0<b_{-}<b_{+}<\infty$, $1<\gamma_{-}<\gamma_{+}\leq 2$, $10<\eta_{-}<\eta_{+}<$ 1. Let $T_{1},T_{2},\ldots$ be a sequence of maps in (3.14), such that the parameters $a_{i},b_{i},\eta_{i},\gamma_{i}$ associated to $T_{i}$ satisfy
@@
Following the representation in [20], we start by describing several partitions associated with nonstationary compositions $T_{n}\circ\cdots\circ T_{1}$ of maps in (3.14). Let $X=[-1,1]$ and, as in Section 3.2.1, denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,+}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively. Here $X_{-}=(-1,0)$ and $X_{+}=(0,1)$. Denote by $\lambda$ the Lebesgue measure on $X$ normalized to probability.
@@
[…]
@@
$G_{\ell}(k)$ =#($=\#\bigl{(}\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr{)},$
@@
$Y^{\prime}_{j}$ =$=\{y\in Y_{k+n-j}:T_{k+n-j,k+\ell-1}(y)\notin Y_{k+\ell}\text{ for all $n-j<\ell\leq n$}\},$
@@
where $\mu_{n}$, $\mu^{\prime}_{n}$ are probability measures with $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n}$ and is $(T_{k,k+n-1})_{*}\mu^{\prime}_{n}$ regular with tail bound $h_{n-N}^{k+N}$ w.r.t. $T_{k+n},T_{k+n+1},\ldots$
@@
[…]
@@
This way, ${\hat{r}}$ is nonincreasing and ${\hat{r}}(1)=1$; for a probability measure, the tail bounds $r$ and ${\hat{r}}$ are equivalent. Similarly define ${\hat{h}}^{k}_{n}$.
@@
To estimate (4.7), we decompose the sum $\sum_{1\leq\ell\leq n+1}H_{\ell}$ intro three parts. To this end, fix $b\in(1/2,1)$ by $b=\bigl{(}1+(1-\theta)^{\frac{1}{2\beta^{\prime}}}\bigr{)}/2$ and let $A\geq 1$ be a large integer whose value will be specified later. Restrict to $\Theta$ sufficiently small so that $01-b-\Theta b>$ 0. Let $C$ denote various constants which depend only on $A,n_{0},\theta,\Theta,C_{h},C_{\beta},\beta,\beta^{\prime}$. In particular, $C$ is independent of $k$.
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 3.3. Grossmann–Horner maps
We follow the presentation in [20]. Let $T:X\to X$ be a map on X=[1,1]X=[-1,1] with graph as in Figure 1, having two surjective and symmetric branches on the subintervals [1,0][-1,0] and $[0,1]$, such that $T|_{X\setminus\{0\}}\in C^{1}$ and $T|_{X\setminus\{0,-1,1\}}\in C^{2}$, defined locally by
@@
Here, $a,b>0$, $\eta>0$ and $\gamma>1$ are parameters, $U_{0}(\gamma,\eta)$ is a small neighborhood of 0, and $U_{1_{-}}(\gamma,\eta),U_{-1_{+}}(\gamma,\eta)$ are small neighborhoods of 1,11,-1, respectively. For $i=1,2$, the derivatives of $T$ are assumed to satisfy
@@
  • $T$ is strictly monotonous and convex on the two intervals [1,0][-1,0] and $[0,1]$, and
@@
Fix $0<a_{-}<a_{+}<\infty$, $0<b_{-}<b_{+}<\infty$, $1<\gamma_{-}<\gamma_{+}\leq 2$, $0<\eta_{-}<\eta_{+}<1$. Let $T_{1},T_{2},\ldots$ be a sequence of maps in (3.14), such that the parameters $a_{i},b_{i},\eta_{i},\gamma_{i}$ associated to $T_{i}$ satisfy
@@
Following the presentation in [20], we start by describing several partitions associated with nonstationary compositions $T_{n}\circ\cdots\circ T_{1}$ of maps in (3.14). Let X=[1,1]X=[-1,1] and, as in Section 3.2.1, denote by $g_{k,-}:(-1,1)\to X_{-}$ and $g_{k,+}:(-1,1)\to X_{+}$ the left and right inverse branches of $T_{k}$, respectively. Here $X_{-}=(-1,0)$ and $X_{+}=(0,1)$. Denote by $\lambda$ the Lebesgue measure on $X$ normalized to probability.
@@
Moreover, $\sup_{k}\int\tau_{k}^{p}\,dm_{k}\leq\sup_{k}\lambda(Y_{k})^{-1}\int\tau_{k}^{p}\,d\lambda<\infty$ holds for some $p>1$ by Proposition 3.18 below, since $\eta_{+}(\gamma_{+}-1)<1$. Now, (NU:5) follows by Proposition 2.1.
@@
$G_{\ell}(k)$ $=\#\bigl(\{n_{j}\>:\>j\in{\mathbb{N}}\}\cap[k,k+\ell-1]\bigr),$
@@
$Y^{\prime}_{j}$ $=\{y\in Y_{k+n-j}:T_{k+n-j,k+\ell-1}(y)\notin Y_{k+\ell}\text{ for all $n-j<\ell\leq n$}\},$
@@
where $\mu_{n}$, $\mu^{\prime}_{n}$ are probability measures with $(T_{k,k+n-1})_{*}\mu_{n}=m_{k+n}$ and $(T_{k,k+n-1})_{*}\mu^{\prime}_{n}$ is regular with tail bound $h_{n-N}^{k+N}$ w.r.t. $T_{k+n},T_{k+n+1},\ldots$
@@
$\sum_{n=1}^{\ell}\xi_{\ell,n}\mu(a_{k,n})$ $=r(\ell)-r(\ell+1)$ for each $\text${ for each $\ell$,
$\sum_{\ell=n}^{\infty}\xi_{\ell,n}$ $=1$ $\text{ for each }n.$
@@
This way, $^{\hat{r}}$ is nonincreasing and ${\hat{r}}(1)=1$; for a probability measure, the tail bounds $r$ and $^{\hat{r}}$ are equivalent. Similarly define ${\hat{h}}^{k}_{n}$.
@@
To estimate (4.7), we decompose the sum $\sum_{1\leq\ell\leq n+1}H_{\ell}$ intro three parts. To this end, fix $b\in(1/2,1)$ by $b=\bigl(1+(1-\theta)^{\frac{1}{2\beta^{\prime}}}\bigr)/2$ and let $A\geq 1$ be a large integer whose value will be specified later. Restrict to $\Theta$ sufficiently small so that $1-b-\Theta b>0$. Let $C$ denote various constants which depend only on $A,n_{0},\theta,\Theta,C_{h},C_{\beta},\beta,\beta^{\prime}$. In particular, $C$ is independent of $k$.
@@
$\leq C_{\beta}^{\prime}C_{2}(\Theta_{k}^{*}k+1)^{\beta^{\prime}}\theta(1-\theta)^{-1}\sum_{j\geq 1}(1-\theta)^{j/2}n^{-\beta^{\prime}}\leq C_{\beta}^{\prime}C_{3}(\Theta_{k}^{*}k+1)^{\beta^{\prime}}n^{-\beta^{\prime}},$
@@
Finally, to see that $\sup_{k}\int\tau_{k}^{p}\,dm_{k}<\infty$ implies (A.1), suppose that $j,k\geq 1$. Since $m_{k}$ is a measure on $Y_{k}$ with $|m_{k}|_{{\mathrm{LL}},k}\leq K_{2}$, arguing as in the proof of Proposition 4.1(a), write
@@
Using $\sup_{k}\int\tau_{k}^{p}\,dm_{k}<\infty$ together with Markov’s inequality, we obtain